Find the equations of the tangents and normal to the following curves:


1. y2 + 8x = 0, parallel to x + y + 4 = 0.

2. x2 = 3y, perpendicular to x – 2y + 7 = 0.

3. x2 + 9y2 = 25, parallel to 4x + 9y + 30 = 0.

4. 25x2 + 4y2 = 100, perpendicular to 8x – 15y + 4 = 0.

5. x2 – y2 = 15, parallel to 4x – y + 20 = 0.



1
Expert's answer
2021-11-29T08:33:33-0500

equation of the tangent:

"y-y_0=f'(x_0)(x-x_0)"

equation of the normal:

"y-y_0=-(x-x_0)\/f'(x_0)"


1.


"y^2+8x=0""2yy'=-8""y'=\dfrac{-4}{y}""slope=m_1=y'=\dfrac{-4}{y}"

Tangent is parallel to "x+y+4=0."



"y=-x-4, slope=m_2=-1"

Then "m_1=m_2=>-\dfrac{4}{y}=-1=>y=4"

Find "x"



"(4)^2+8x=0=>x=-2"

Point "(-2, 4)"

Tangent line is



"y-4=-(x-(-2))"

The equation of the tangent line in slope-intercept form



"y=-x+2"



Find the equation of the normal at the point "(-2, 4)"



"slope=m_3=-\dfrac{1}{m_1}=1""y-4=(x-(-2))"

The equation of the normal line in slope-intercept form



"y=x+6"

2.


"x^2=3y""2x=3y'""y'=\dfrac{2x}{3}""slope=m_1=y'=\dfrac{2x}{3}"

Tangent is perpendicular to "x-2y+7=0."

"y=\dfrac{1}{2}x+\dfrac{7}{2}, slope=m_2=\dfrac{1}{2}"

Then "m_1=-\dfrac{1}{m_2}=>\dfrac{2x}{3}=-2=>x=-3"

Find "y"

"(-3)^2=3y=>y=3"

Point "(-3, 3)"

Tangent line is



"y-3=-2(x-(-3))"

The equation of the tangent line in slope-intercept form



"y=-2x-3"



Find the equation of the normal at the point "(-3, 3)"

"slope=m_3=-\dfrac{1}{m_1}=\dfrac{1}{2}""y-3=\dfrac{1}{2}(x-(-3))"

The equation of the normal line in slope-intercept form



"y=\dfrac{1}{2}x+\dfrac{9}{2}"



3.


"x^2+9y^2=25""2x+18yy'=0""y'=\dfrac{-x}{9y}""slope=m_1=y'=-\dfrac{x}{9y}"

Tangent is parallel to "4x + 9y + 30 = 0."



"y=-\dfrac{4}{9}x-\dfrac{10}{3}, slope=m_2=-\dfrac{4}{9}"

Then "m_1=m_2=>-\dfrac{x}{9y}=-\dfrac{4}{9}=>x=4y"

Find "y"



"(4y)^2+9y^2=25=>y=\pm1"

Point "(-4, -1)," Point "(4, 1)."


Point "(-4, -1)"

Tangent line is



"y-(-1)=-\dfrac{4}{9}(x-(-4))"

The equation of the tangent line in slope-intercept form



"y=-\dfrac{4}{9}x-\dfrac{25}{9}"



Find the equation of the normal at the point "(-4, -1)"



"slope=m_3=-\dfrac{1}{m_1}=\dfrac{9}{4}""y-(-1)=\dfrac{9}{4}(x-(-4))"

The equation of the normal line in slope-intercept form



"y=\dfrac{9}{4}x+8"

Point "(4, 1)"

Tangent line is



"y-1=-\dfrac{4}{9}(x-4)"

The equation of the tangent line in slope-intercept form



"y=-\dfrac{4}{9}x+\dfrac{25}{9}"



Find the equation of the normal at the point "(4, 1)"



"slope=m_3=-\dfrac{1}{m_1}=\dfrac{9}{4}""y-1=\dfrac{9}{4}(x-4)"

The equation of the normal line in slope-intercept form



"y=\dfrac{9}{4}x-8"

4.


"25x^2+4y^2=100""50x+8yy'=0""y'=\dfrac{-25x}{4y}""slope=m_1=y'=-\dfrac{25x}{4y}"

Tangent is perpendicular to "8x \u2013 15y + 4 = 0."



"y=\dfrac{8}{15}x+\dfrac{4}{15}, slope=m_2=\dfrac{8}{15}"

Then "m_1=-\dfrac{1}{m_2}=>-\dfrac{25x}{4y}=-\dfrac{15}{8}=>x=\dfrac{3}{10}y"

Find "y"



"25(\dfrac{3}{10}y)^2+4y^2=100=>y=\pm4"

Point "(-\dfrac{6}{5}, -4)," Point "(\dfrac{6}{5}, 4)."


Point "(-\dfrac{6}{5}, -4)"

Tangent line is



"y-(-4)=-\dfrac{15}{8}(x-(-\dfrac{6}{5}))"

The equation of the tangent line in slope-intercept form



"y=-\dfrac{15}{8}x-\dfrac{25}{4}"



Find the equation of the normal at the point "(-\dfrac{6}{5}, -4)"

"slope=m_3=-\dfrac{1}{m_1}=\dfrac{8}{15}""y-(-4)=\dfrac{8}{15}(x-(-\dfrac{6}{5}))"

The equation of the normal line in slope-intercept form



"y=\dfrac{8}{15}x-\dfrac{84}{25}"

Point "(\dfrac{6}{5}, 4)"

Tangent line is



"y-4=-\dfrac{15}{8}(x-\dfrac{6}{5})"

The equation of the tangent line in slope-intercept form



"y=-\dfrac{15}{8}x+\dfrac{25}{4}"



Find the equation of the normal at the point "(\dfrac{6}{5}, 4)"



"slope=m_3=-\dfrac{1}{m_1}=\dfrac{8}{15}""y-4=\dfrac{8}{15}(x-\dfrac{6}{5})"

The equation of the normal line in slope-intercept form



"y=\dfrac{8}{15}x+\dfrac{84}{25}"

5.


"x^2-y^2=15""2x-2yy'=0""y'=\dfrac{x}{y}""slope=m_1=y'=\dfrac{x}{y}"

Tangent is parallel to "4x \u2013 y + 20 = 0."



"y=4x+20, slope=m_2=4"

Then "m_1=m_2=>\dfrac{x}{y}=4=>x=4y"

Find "y"



"(4y)^2-y^2=15=>y=\pm1"

Point "(-4, -1)," Point "(4, 1)."


Point "(-4, -1)"

Tangent line is



"y-(-1)=4(x-(-4))"

The equation of the tangent line in slope-intercept form



"y=4x+15"



Find the equation of the normal at the point "(-4, -1)"



"slope=m_3=-\dfrac{1}{m_1}=-\dfrac{1}{4}""y-(-1)=-\dfrac{1}{4}(x-(-4))"

The equation of the normal line in slope-intercept form



"y=-\dfrac{1}{4}x-2"

Point "(4, 1)"

Tangent line is



"y-1=4(x-4)"

The equation of the tangent line in slope-intercept form



"y=4x-15"



Find the equation of the normal at the point "(4, 1)"



"slope=m_3=-\dfrac{1}{m_1}=-\dfrac{1}{4}""y-1=-\dfrac{1}{4}(x-4)"

The equation of the normal line in slope-intercept form



"y=-\dfrac{1}{4}x+2"